### v3

@@ 3.3 Detecting Invalid Propagations in the Presolving Library PaPILO
PaPILO chooses to apply Probing [27, 17] to the variable for (LABEL:eq::3662). Probing tentatively fixes binary variables to 0 and 1, applies constraint propagation [27] and tries to conclude implication based on the outcomes of the propagation. By fixing to 0, the problem correctly propagates to infeasibility. If the variable is fixed to , propagation tries to find new bounds for the variable by resolving the constraints by :
@@
In this case, is first ceiled to 0 against .
@@ 3.4 Identifying a Numerically Critical Constraint Normalization
Since PaPILO is not called in the original instance fixing the bug in PaPILO does not fix the solve status of the original instance. Therefore, re-running MIP-DD with the fixed PaPILO and suppressed primal fails results in a new reduced problem
@@
In this problem, the second constraint leads to numerical difficulties. During presolving, all variables except are either fixed () or substituted resulting in the following problem
SCIP normalizes the remaining constraint by dividing it by leading to a right hand side that is smaller than the feasible tolerance :
@@ 3.5 Fixing the Solution Count
For example, in the counting mode of SCIP, if a problem is already solved during presolving, a solution count of zero was claimed, even though a feasible solution was provided when using the regular optimization mode.
@@ 4 Conclusion and Outlook
To conclude, let us revisit different benefits of delta debugging during the development process of an MIP solver and give an outlook on possible enhancements for the future. First, as shown in Section 3, MIP-DD is able to fulfill its general purpose of reducing large instances to tiny instances while maintaining the error. This spares a lot of time during debugging by reducing the hurdles for investigating issues because all relevant calculations done by the solver on the reduced instance can usually be traced and checked manually. Further, MIP-DD can be applied even if other techniques such as providing a debug solution fail.
Second, in the context of collaborations with industry, it regularly happens that instances cannot be shared due to the General Data Protection Regulation or other confidentiality reasons. In this case, a reduced instance with renamed variable and constraints names obtained by delta debugging may barely contain any sensitive information anymore about the original instances, making the reduced instance safe for customers to share with developers.
Third, the MIP-DD itself does not detect an error, rather is the definition of an error implemented in the solver-API. By this design choice, MIP-DD can not only be applied to detect bugs caused by primal or dual infeasibility but rather to analyze any unwanted behavior specific to a certain solver. As one example, the upcoming MIP-DD 2.0 will contain, along with improvements for automatic limit settings, adaptive batch numbers, and extended arithmetical options, an interface to the numerically exact version of SCIP [28, 29, 21] and the verification software VIPR for checking validity of certificates [30]. The generation of an invalid certificate can then be chosen as another possible error to be reproduced, and reducing the instance size goes hand in hand with producing a small, invalid VIPR certificate.
In an even broader context, MIP-DD could also be utilized to track down performance bottlenecks for example by letting the solver return an error at a certain tree depth which should not be reached on a given instance. This way, issues with exploding tree sizes on instances turning out to be simple could be investigated manually, which contributes to creating more performant solvers in itself.
@@ Acknowledgements.
Furthermore, we wish to thank Leon Eifler for helping to implement the interface to the exact version of SCIP and SoPlex, Felipe Serrano for his careful proofreading, and Armin Biere for sharing his experience and perspective on delta debugging in the SAT solver world.

### v5

@@ 3.3 Detecting Invalid Propagations in the Presolving Library PaPILO
PaPILO chooses to apply Probing [29, 19] to the variable . Probing tentatively fixes binary variables to 0 and 1, applies constraint propagation [29] and tries to conclude implication based on the outcomes of the propagation. By fixing to 0, the problem correctly propagates to infeasibility. If the variable is fixed to , propagation tries to find new bounds for the variable by resolving the constraints by :
@@
In this case, is first ceiled to 0 against the feasibility tolerance .
@@ 3.4 Identifying a Numerically Critical Constraint Normalization
Since PaPILO is not called in the original instance fixing the bug in PaPILO does not fix the incorrect solver status of the original instance discussed in Section 3.3. Therefore, re-running MIP-DD with the fixed PaPILO and suppressed primal fails results in a new reduced problem
@@
In this problem, the second constraint leads to numerical difficulties. During presolving, all variables except are either fixed ( and ) or substituted resulting in the following problem
SCIP normalizes the remaining constraint by dividing it by leading to a right hand side that is smaller than the feasibility tolerance :
@@ 3.5 Fixing the Solution Count
For example, SCIP incorporates a counting mode, in which the number of all feasible value assignments on integral variables is counted333https://www.scipopt.org/doc/html/COUNTER.php.
If a problem is already solved during presolving, a solution count of zero was claimed, even though a feasible solution was provided when using the regular optimization mode.
@@ 4 Conclusion and Outlook
To conclude, let us revisit the different benefits of delta debugging during the development process of an MIP solver and give an outlook on possible enhancements for the future. First, as shown in Section 3, MIP-DD is able to fulfill its general purpose of reducing the instance size while still providing incorrect output. This spares a lot of time during debugging by reducing the hurdles for investigating issues because all relevant calculations done by the solver on the reduced instance can usually be traced and checked manually. Further, MIP-DD can be applied even if other techniques such as providing a debug solution fail.
Second, in the context of collaborations with industry, it regularly happens that instances cannot be shared due to the General Data Protection Regulation [30] or other confidentiality reasons. In this case, a reduced instance with renamed variable and constraint names obtained by delta debugging may barely contain any sensitive information anymore about the original instances, making the reduced instance safe for customers to share with developers.
Third, MIP-DD itself does not prescribe what counts as an error. Instead, the definition of an error is provided by the implementation of the solver API. Hence, MIP-DD can not only be applied to detect bugs caused by primal or dual infeasibility but can analyze any unwanted behavior specific to a certain solver. As one example, the MIP-DD 2.0 will contain, along with improvements for automatic limit settings, adaptive batch numbers, and extended arithmetical options, an experimental interface to the numerically exact version of SCIP [31, 32, 23] and the verification software VIPR for checking validity of certificates [33]. This will make it possible to identify the generation of an invalid certificate as another type of failure. In this context, reducing the size of problem instances is helpful because it typically goes hand in hand with reducing the size of invalid certificates, which therefore may become easier to analyze.
In an even broader context, MIP-DD could also be utilized to track down performance bottlenecks, e.g. by letting the solver return an error at a certain tree depth which should not be reached on a given instance. This way, issues with exploding tree sizes on instances turning out to be simple could be investigated manually, which contributes to creating more performant solvers in itself.
@@ Acknowledgements.
Furthermore, we wish to thank Leon Eifler for helping to implement the interface to the exact version of SCIP and SoPlex, Mark Turner and Felipe Serrano for their careful proofreading, and Armin Biere for sharing his experience and perspective on delta debugging in the SAT solver world.
Finally, we would like to thank the associate editor and two anonymous reviewers for their insightful comments and suggestions, which have contributed to improving the quality of this work.